OpenStudy (anonymous):

HELP!!! I will give medal to anyone who can explain how they got their answer. Evaluate if it exists:

OpenStudy (anonymous):

\[\int\limits_{0}^{\ln(2)}x ^{-2}e ^{-1/x}dx\]

OpenStudy (anonymous):

u sub. let u = 1/x then du = -1/x^2=-x^-2

OpenStudy (anonymous):

Thanks again can you answer one more question?

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ 1 }{ \sqrt{7+6x-x ^{2}} }\]There is no bounds just integral

OpenStudy (anonymous):

is that everything?

OpenStudy (anonymous):

Yes no bounds just integral

OpenStudy (anonymous):

\[\int\frac{dx}{7+6x-x^{2}}\] like this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

No the denominator is square-rooted

OpenStudy (anonymous):

oops... I meant \[\int\frac{dx}{\sqrt{7+6x-x^{2}}}\] like this?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

you need to complete the square of the polynomial inside the square root, i.e., complete the square on 7+6x_x^2. then you it should be an inverse trig. function (once integrated).

OpenStudy (anonymous):

When you mean complete the square, do you mean factor?

OpenStudy (anonymous):

no, i mean complete the square. look at this website http://www.math.hmc.edu/calculus/tutorials/complete_sq/

OpenStudy (anonymous):

I got x^2+6x+9=2 is that right?

OpenStudy (anonymous):

it should be 16-(x-3)^2 = 16 - x^2 +6x -9 = 7+6x-x^2

OpenStudy (anonymous):

Oh ok so is the trig identity the sine inverse one?

OpenStudy (anonymous):

I think so... it's been a while

OpenStudy (anonymous):

Thanks for the help.